OpenStudy (anonymous):

If u = (6 , 1 , 8) is orthogonal to u = (k , -4 , 5) determine the value of k. Please help! Thanks :)

OpenStudy (ajprincess):

u is orthogonal to u means u.u=0. so 6k-4+8*5=0 6k-4+40=0 6k-36=0 6k=36 k=36/6 k=6

OpenStudy (anonymous):

Ohhh, I seee... THANKS! :)

OpenStudy (ajprincess):

yw.:)